# Revisiting the question of what instantaneous normal modes tell us about liquid dynamics

### v1

@@ IV.1 Fraction of unstable modes
In all liquids, the fraction of unstable modes increases with temperature. A larger kinetic energy corresponds to a larger probability of exploring regions of the potential landscape with locally negative curvature, such as saddle points or maxima, moving the dynamics away from the bottom of the local minima. Despite this might seem natural from a physical perspective, how $f_{u}$ increases with temperature is much less trivial. In order to provide valuable information regarding the temperature trend of $f_{u}$, we have collected all our results in Fig.7. In panel a, we show $f_{u}$ for Ar, Xe, Ga and Pb. For a better comparison, we normalized the data using the melting temperature of each of these liquids $T_{m}$ and the corresponding value of $f_{u}$ at melting. We observe that $f_{u}(T)$ is well described in all cases by a power-law function $f_{u}(T)\sim T^{\alpha}$, where nevertheless $\alpha$ depends on the specific liquid chosen. More precisely we find that $\alpha$ is $1.09,1.06,0.84,0.71$ for Ar, Xe, Ga and Pb respectively. For N2 and CS2 we do not have data up to the melting temperature and we therefore decided to normalize the data by the minimum temperature available, $T_{0}$, and show the corresponding results in a separate panel, panel b in Fig.7. There, we find again a trend consistent with a single power-law, where the power $\alpha$ is $1.14$ for N2 and $0.88$ for CS2.
It is noteworthy that the power-law $\alpha$ is close to $1$ for all the liquids considered. This is in good agreement with the predictions from the random energy model (see e.g. Keyes (2005)), according to which $f_{u}(T)=\frac{T}{\delta 2\pi}e^{-\delta^{2}/(4T^{2})}$ where $\delta$ is an $\mathcal{O}(1)$ parameter describing the width of the energy distribution. We have tried to fit the data with the expression from the random energy model with a single free parameter $\delta$ and found that it does not capture correctly the trend of the numerical data.
@@ IV.2 Low frequency linear scaling
This behavior can be derived using several theoretical arguments Xu and Stratt (1989); Wu and Loring (1993); Wan and Stratt (1994); Zaccone and Baggioli (2021); Schirmacher et al. (2022); Li and Keyes (1999); Zürcher and Keyes (1997); Wu and Loring (1992) and it is directly related to the presence of unstable modes with imaginary frequency. In all the systems considered in this manuscript, that include Ar, Xe, Ga, Pb, N2 and CS2, the INM DOS follow the linear form presented in Eq.(13). In general terms, much less is known about the temperature dependence of the parameter $a(T)$ that characterizes the slope of this linear trend.
In a series of works Keyes (1994b); Li and Keyes (1999) (see Keyes (2005) for a review), Keyes and collaborators proposed a closed-form formula for this parameter that, under some reasonable approximations, takes the following form
@@
Here, $f_{u}^{\text{max}}$ is the maximum value for the fraction of unstable modes, and depends on the details of the potential landscape. On the other hand, $\langle E\rangle$ is an average energy scale that is supposedly related to hopping processes and the anharmonic nature of the potential landscape. As seen in the previous Section, the fraction of unstable modes grows with temperature. This implies that the first term in Eq.(14) is a term that decreases with temperature. The other contribution to Eq.(14) is an exponential growth of the scaling upon increasing the temperature, whose growth rate is controlled by the energy scale $\langle E\rangle$. Because of the competition between these two terms, in general Eq.(14) predicts that $a(T)$ grows/decreases with temperature depending on which of the terms therein dominates. These two behaviors have been both observed in simulations. For a Lennard Jones liquid Keyes (1994b), the energy scale $\langle E\rangle$ is very low; hence, the first term in Eq.(14) dominates the behavior of $a(T)$ that decreases with temperature. The same is observed for bulk polymers Zhang et al. (2019). On the contrary, for many other systems, including CS2 Li and Keyes (1999) and water Jin et al. (2024), $a(T)$ grows with temperature following a form compatible with the exponential term in Eq.(14).
Figure 8: The INM slope extracted from the experimental data using Eq.(13) as a function of the inverse temperature $1/T$. The red dashed line in panels (a)-(d) is the result of an exponential fit $a(t)\propto\exp(-\langle E\rangle/k_{B}T)$. For all systems, the data points are above the melting temperature.
@@
We notice that the reported value of $\langle E\rangle\approx 13\,\text{meV}$ for water Jin et al. (2024) is of the same order of magnitude of the values found above. As we will see in the next Section, these activation energies are consistently smaller than those governing the Arrhenius behavior of the corresponding diffusion constants. For example, for $N_{2}$ the activation energy characterizing the slope $a(T)$ is about $20$ times smaller than the one describing $D(T)$. This implies that, despite the same functional form, it is very unlikely that the microscopic origin of it is the same for the slope $a(T)$ and the diffusion constant $D(T)$. Nevertheless, we notice that all the normal modes analysis, including therefore the determination of $a(T)$, is performed using the harmonic approximation. On the contrary, the determination of $D(T)$ does not rely on this approximation. It is therefore conceivable (similar arguments have been made in Jin et al. (2024) by comparing the INM slope and the experimental slope) that the strong liquid anharmonicities are the responsible for this difference.
@@ IV.3 Self-diffusion constant
One of the initial motivations to introduce and study instantaneous normal modes was the idea of reconstructing the liquid self-diffusion constant in terms of liquid normal modes Zwanzig (1983). Keyes and collaborators have proposed that the self-diffusion constant in liquids could be related to the unstable instantaneous normal modes Madan et al. (1990b), and that a somehow simple relation between $D$ and $f_{u}$ could exist,
@@
Despite INM describes diffusion in liquids well Keyes (1997b, 1994b); Madan et al. (1990b); Keyes (1995), several open questions remain. First, it is by now well-known that the total fraction of unstable modes cannot be a direct source for self-diffusion since $D=0$ in solids but $f_{u}$ is not. As already mentioned above, this has motivated a large number of studies to find out which subset of $f_{u}$ is actually responsible for self-diffusion (e.g., La Nave et al. (2001); Sciortino and Tartaglia (1997) and discussions therein). Moreover, in several instances (e.g., CS2) it has been shown that Eq.(16) does not hold and a “stretched” version of Eq.(16) has to be considered (see e.g. Li and Keyes (1999)),
@@
We have computed the temperature dependent self-diffusion constant $D(T)$ for our six liquid systems. The behavior of $D$ as a function of temperature is presented in Arrhenius form in Appendix 4. All data are consistent with an Arrhenius behavior,
@@
In order to investigate further the relation between the self-diffusion constant $D$ and the fraction of unstable modes, in Fig.9 we have plotted $D$ as a function of $f_{u}$. In order to compare the different systems, we have normalized both quantities by subtracting their value at the melting temperature $T_{m}$ (for Ar, Xe, Ga and Pb) or at the minimum temperature available $T_{0}$ (for N2 and CS2). For all the systems considered, we observe a power law scaling, with the power $\beta$ ranging from $1.15$ in Xe to $1.96$ in CS2. We notice that most of the powers obtained by fitting are not far from the value of $1$, that would be consistent with Eq.(16) and the analysis in Li and Keyes (1999). Nevertheless, we also observe important deviations and we do not find any universal scaling.
@@ IV.4 Low temperature solid-like limit
Let us consider a liquid. Its INM DOS is linear in the frequency. On the other hand, Debye taught us that for a zero temperature ideal crystal, the corresponding vibrational density of states has a quadratic behavior, known as Debye’s law Kittel (2021). The contrast between these two opposite trends has been observed experimentally in selenium Phillips et al. (1989), heavy water Dawidowski et al. (2000) and water Jin et al. (2024). Less is known about how the INM DOS interpolates between these two behaviors by lowering the temperature and approaching the ideal crystalline phase. As already mentioned, the presence of unstable modes (no matter how few) implies that a linear scaling appears in the INM DOS. Therefore, as soon as unstable mode are present in the spectrum and $f_{u}\neq 0$, the INM DOS will be linear at low enough frequencies. By decreasing temperature, the fraction of unstable modes diminishes and the linear scaling is expected to get confined to very low frequency and eventually give room to the Debye scaling, that is allowed to appear only when $f_{u}=0$.
Figure 10: (a) The experimental density of states for liquid gallium at $500$ K reproduced with permission from Ref.Stamper et al. (2022). (b) The INM DOS of Gallium simulated at different temperatures. The first-order melting temperature is between 320 K and 340 K. (c) The density of state function $S(\omega)$ derived from the Fourier transform of the VACF for the same simulated Gallium system.
In order to investigate this point, we have chosen Gallium as a benchmark example. As a reference, in Fig.10, we present the experimental data for the density of states at $500$ K from Ref.Stamper et al. (2022) together with our simulation results for the INM density of states $\rho(\omega)$ and the density of states function $S(\omega)$. Among the various features emerging from this comparison, we notice that at low temperature (e.g., $T=300$ K), $S(\omega)$ shows already a clear Debye scaling at low frequency and a vanishing diffusion constant ($S(0)=0$). On the contrary, the stable branche of the INM DOS is still strongly linear at low frequency, consistent with a large number of unstable mode still present at such a temperature.
@@
In order to check this point further, in panel (b) of Fig.11, we have zoomed the stable part of the INM spectrum of Gallium at $T=1$ K and presented it in a log-log scale to highlight possible power-law scalings. The data present a clear linear scaling up to approximately $4.5$ meV, that is a consequence of the remaining unstable modes (and anharmonicities). On the other hand, around $\approx 4.5$ meV we observe a clear crossover to a larger power that is consistent with the Debye scaling $\rho(\omega)\propto\omega^{2}$ and that persists to $\approx 8$ meV. This reveals the approach of the solid-like behavior to low frequency by decreasing the temperature. By decreasing further $T$, one expects that the crossover frequency will move towards the origin, and finally vanish at $T=0$ when $f_{u}=0$. We notice that an alternative method to observed this trend, and the emergence of the Debye scaling, consists in nano-confining the liquid, as experimentally performed in Yu et al. (2024). We also notice that the presence of unstable modes in finite temperature (amorphous) solids has been verified experimentally, see for example Vaibhav et al. (2024).
@@ IV.5 Zero eigenvalue cusp singularity
Following previous observations Sastry et al. (2001); Taraskin and Elliott (2002), Refs. Schirmacher et al. (2022); Mossa et al. (2023) have recently pointed out that the frequency representation of the INM density of states $\rho(\omega)$ might hide important information that are contained in the corresponding eigenvalue distribution $\upsilon(\lambda)$. Their arguments are supported by numerical simulations in a soft-sphere liquid model (data reproduced in panel (a) of Fig.14). Since $\lambda=\omega^{2}$, the eigenvalue distribution and the INM DOS are simply related via the Jacobian,
@@
Mathematically, it appears therefore unreasonable that one of the two quantities contain more information than the other. Nevertheless, in this Section we entertain this possibility and analyze in detail the arguments of Schirmacher et al. (2022); Mossa et al. (2023) in a large set of liquid models. First, we would like to point out that the statement that the INM density of states is universally linear at low frequencies is equivalent to the statement that $\rho(0)\neq 0$, that is certainly true. In other words, the presence of a cusp at $\lambda=0$ is not incompatible with a low-frequency form $\rho(\omega)=a(T)\omega+\dots$, observed in many instances.
@@
In order to characterize these features in a more quantitative way, we resort to the empirical function proposed in Mossa et al. (2023),
@@
Figure 12: The eigenvalue distribution $\upsilon(\lambda)$ for different temperatures. From top left to bottom right: Ar (40 bar), Xe ($1.01$ bar), Pb($1.01$ bar), Ga($1.01$ bar), N2 ($70.92$ bar) and CS2 ($1.01$ bar).
@@
In panel (a) of Fig.13, we show the temperature behavior of the fitted stable (red) and unstable (black) slopes as a function of temperature for the case of CS2. The slopes are extracted from the numerical data using Eq.(21). Similar results are obtained for the different liquids shown in Fig.12. We observe that the positive slope of the unstable branch grows with temperature while the negative one, corresponding to the unstable part of the spectrum, moves towards larger negative values by increasing $T$. This trend is opposite of what reported in Schirmacher et al. (2022); Mossa et al. (2023) for the soft sphere model. The reason for this difference is ultimately the opposite behavior of $\upsilon(0)$ with temperature in the soft sphere model. In particular, as evident from Fig.1 in Schirmacher et al. (2022), $\upsilon(0)$ in the soft sphere model decreases by increasing temperature. In our language, this corresponds to an INM slope $a(T)$ decreasing with temperature, as for LJ liquids, but opposite to our more realistic liquid systems.
@@
This contrast is then reflected in the temperature dependence of the difference of the slopes, that is presented in panel (b) of Fig.13. There, we observe that $\Delta_{\text{slope}}$ increases with temperature, following a trend approximately consistent with a linear power-law scaling. This is again very different from the $T^{-2/3}$ behavior observed in the soft sphere model Schirmacher et al. (2022); Mossa et al. (2023), not only for the precise power involved in the scaling, but even in the qualitative trend as well. Our findings imply that the $T^{-2/3}$ law found in Schirmacher et al. (2022); Mossa et al. (2023) is not universal and its value and significance should be revisited.
@@
Finally, we would like emphasize that the theoretical model presented in Schirmacher et al. (2022); Mossa et al. (2023) is not the only one able to predict the presence of a cusp-like singularity in the eigenvalue distribution of liquids. Indeed, a much older model known as the “soft potential model of liquids” (SPM) Zürcher and Keyes (1997); Xu et al. (2023) already captures such a feature. We provide a direct proof of this statement in panel b of Fig.14. The presence of a cusp singularity at $\lambda=0$ is evident. For completeness, in the inset we also show the INM DOS that shows the aforementioned linear in frequency regime despite the cusp.
@@
Figure 14: (a) The eigenvalue distribution $\upsilon(\lambda)$ for the soft-sphere liquid model presented in Mossa et al. (2023) and the corresponding fraction of unstable mode $f_{u}$ as a function of temperature in reduced LJ units. $T_{g}$ is the glass temperature of the corresponding system. (b) The eigenvalue distribution $\upsilon(\lambda)$ obtained from the theoretical soft potential model of liquids Zürcher and Keyes (1997); Xu et al. (2023) by converting the calculated $\rho(\omega)$ to $\upsilon(\lambda)$ with the Jacobian relation. The inset shows the original INM DOS $\rho(\omega)$ with the dashed black lines indicating the universal linear regime $\rho(\omega)\propto\omega$ which emerges clearly despite the presence of the cusp-like singularity in $\upsilon(\lambda)$.
@@
In this direction, it would be interesting to understand further the connection between the theory presented in Schirmacher et al. (2022); Mossa et al. (2023) and the older SPM Zürcher and Keyes (1997); Xu et al. (2023). We notice that the two frameworks start from completely different assumptions. The heterogeneous elasticity theory of Schirmacher et al. (2022); Mossa et al. (2023) follows a harmonic description, whose validity is arguable in the liquid phase, especially at large temperatures. On the other hand, the SPM starts from a anharmonic picture of the dynamics that involves additional low-energy modes. Also in this case, the nature and origin of these modes in liquids remains an open question. String-like excitations Donati et al. (1998) can represent a candidate for these anharmonic excitations.

### v2

@@ IV.1 Fraction of unstable modes
In all liquids, the fraction of unstable modes increases with temperature. A larger kinetic energy corresponds to a larger probability of exploring regions of the potential landscape with locally negative curvature, such as saddle points or maxima, moving the dynamics away from the bottom of the local minima. Despite this might seem natural from a physical perspective, how $f_{u}$ increases with temperature is much less trivial. In order to provide valuable information regarding the temperature trend of $f_{u}$, we have collected all our results in Fig.7. In panel a, we show $f_{u}$ for Ar, Xe, Ga and Pb. For a better comparison, we normalized the data using the melting temperature of each of these liquids $T_{m}$ and the corresponding value of $f_{u}$ at melting. We observe that $f_{u}(T)$ is well described in all cases by a power-law function $f_{u}(T)-f_{u}(T_{m})\sim\left(T-T_{m}\right)^{\alpha}$, where nevertheless $\alpha$ depends on the specific liquid chosen. More precisely we find that $\alpha$ is $1.09,1.06,0.84,0.71$ for Ar, Xe, Ga and Pb respectively. For N2 and CS2 we do not have data up to the melting temperature and we therefore decided to normalize the data by the minimum temperature available, $T_{0}$, and show the corresponding results in a separate panel, panel b in Fig.7. There, we find again a trend consistent with a single power-law, where the power $\alpha$ is $1.14$ for N2 and $0.88$ for CS2.
It is noteworthy that the power-law $\alpha$ is close to $1$ for all the liquids considered. This is in good agreement with the predictions from the random energy model (see e.g. [121]), according to which $f_{u}(T)=\frac{T}{\delta 2\pi}e^{-\delta^{2}/(4T^{2})}$ where $\delta$ is an $\mathcal{O}(1)$ parameter describing the width of the energy distribution. We have tried to fit the data with the expression from the random energy model with a single free parameter $\delta$ and found that it does not capture correctly the trend of the numerical data.
@@ IV.2 Low frequency linear scaling
This behavior can be derived using several theoretical arguments [32, 33, 34, 35, 36, 37, 30, 38] and it is directly related to the presence of unstable modes with imaginary frequency. In all the systems considered in this manuscript, that include Ar, Xe, Ga, Pb, N2 and CS2, the INM DOS follow the linear form presented in Eq.(14). In general terms, much less is known about the temperature dependence of the parameter $a(T)$ that characterizes the slope of this linear trend.
In a series of works [19, 37] (see [121] for a review), Keyes and collaborators proposed a closed-form formula for this parameter that, under some reasonable approximations, takes the following form
@@
Here, $f_{u}^{\text{max}}$ is the maximum value for the fraction of unstable modes, and depends on the details of the potential landscape. On the other hand, $\langle E\rangle$ is an average energy scale that is supposedly related to hopping processes and the anharmonic nature of the potential landscape. As seen in the previous Section, the fraction of unstable modes grows with temperature. This implies that the first term in Eq.(15) is a term that decreases with temperature. The other contribution to Eq.(15) is an exponential growth of the scaling upon increasing the temperature, whose growth rate is controlled by the energy scale $\langle E\rangle$. Because of the competition between these two terms, in general Eq.(15) predicts that $a(T)$ grows/decreases with temperature depending on which of the terms therein dominates. These two behaviors have been both observed in simulations. For a Lennard Jones liquid [19], the energy scale $\langle E\rangle$ is very low; hence, the first term in Eq.(15) dominates the behavior of $a(T)$ that decreases with temperature. The same is observed for bulk polymers [65]. On the contrary, for many other systems, including CS2 [37] and water [53], $a(T)$ grows with temperature following a form compatible with the exponential term in Eq.(15).
We notice that a slope of the form $a(T)\propto\exp\left(\langle E\rangle/k_{B}T\right)$ could be also explained from gapped shear waves ($k$-gap theory [110]). Indeed, as pointed out in [113, 114], gapped shear waves would contribute a term of the form $g(\omega)\propto\omega/\tau_{m}$. A relaxation time $\tau_{m}\propto\exp\left(\langle E\rangle/k_{B}T\right)$, that decreases exponentially with temperature, would immediately imply such a temperature dependence for $a(T)$. Nevertheless, this argument would not explain the scenarios (e.g., LJ liquids [19]) where $a(T)$ decreases with temperature.
Figure 8: The INM slope extracted from the experimental data using Eq.(14) as a function of the inverse temperature $1/T$ for Ar (a), Xe (b), N2 (c), CS2 (d), Ga (e) and Pb (f). The red dashed line in panels (a)-(d) is the result of an exponential fit $a(t)\propto\exp(-\langle E\rangle/k_{B}T)$. For all systems, the data points are above the melting temperature.
@@
We notice that the reported value of $\langle E\rangle\approx 13\,\text{meV}$ for water [53] is of the same order of magnitude of the values found above. As we will see in the next Section, these activation energies are consistently smaller than those governing the Arrhenius behavior of the corresponding diffusion constants. For example, for $N_{2}$ the activation energy characterizing the slope $a(T)$ is about $20$ times smaller than the one describing $D(T)$. This implies that, despite the same functional form, it is very unlikely that the microscopic origin of it is the same for the slope $a(T)$ and the diffusion constant $D(T)$. Nevertheless, we notice that all the normal modes analysis, including therefore the determination of $a(T)$, is performed using the harmonic approximation. On the contrary, the determination of $D(T)$ does not rely on this approximation. It is therefore conceivable (similar arguments have been made in [53] by comparing the INM slope and the experimental slope) that the strong liquid anharmonicities are the responsible for this difference.
@@ IV.3 Self-diffusion constant
One of the initial motivations to introduce and study instantaneous normal modes was the idea of reconstructing the liquid self-diffusion constant in terms of liquid normal modes [49]. Keyes and collaborators have proposed that the self-diffusion constant in liquids could be related to the unstable instantaneous normal modes [26], and that a somehow simple relation between $D$ and $f_{u}$ could exist,
@@
Despite INM describes diffusion in liquids well [12, 19, 26, 20], several open questions remain. First, it is by now well-known that the total fraction of unstable modes cannot be a direct source for self-diffusion since $D=0$ in solids but $f_{u}$ is not. As already mentioned above, this has motivated a large number of studies to find out which subset of $f_{u}$ is actually responsible for self-diffusion (e.g., [108, 106] and discussions therein). Moreover, in several instances (e.g., CS2) it has been shown that Eq.(17) does not hold and a “stretched” version of Eq.(17) has to be considered (see e.g. [37]),
@@
We have computed the temperature dependent self-diffusion constant $D(T)$ for our six liquid systems. The behavior of $D$ as a function of temperature is presented in Arrhenius form in .5. All data are consistent with an Arrhenius behavior,
@@
In order to investigate further the relation between the self-diffusion constant $D$ and the fraction of unstable modes, in Fig.9 we have plotted $D$ as a function of $f_{u}$. In order to compare the different systems, we have normalized both quantities by subtracting their value at the melting temperature $T_{m}$ (for Ar, Xe, Ga and Pb) or at the minimum temperature available $T_{0}$ (for N2 and CS2). For all the systems considered, we observe a power law scaling, with the power $\beta$ ranging from $1.15$ in Xe to $1.96$ in CS2. We notice that most of the powers obtained by fitting are not far from the value of $1$, that would be consistent with Eq.(17) and the analysis in [37]. Nevertheless, we also observe important deviations and we do not find any universal scaling.
@@ IV.4 Low temperature solid-like limit
Let us consider a liquid. Its INM DOS is linear in the frequency. On the other hand, Debye taught us that for an ideal crystal, the corresponding vibrational density of states has a quadratic behavior, known as Debye’s law [122]. The contrast between these two opposite trends has been observed experimentally in selenium [50], heavy water [51] and water [53]. Less is known about how the INM DOS interpolates between these two behaviors by lowering the temperature and approaching the ideal crystalline phase (see nevertheless [42] where this transition is observed when only pure translation INM are considered). As already mentioned, the presence of unstable modes (no matter how few) implies that a linear scaling appears in the INM DOS. Therefore, as soon as unstable mode are present in the spectrum and $f_{u}\neq 0$, the INM DOS will be linear at low enough frequencies. By decreasing temperature, the fraction of unstable modes diminishes and the linear scaling is expected to get confined to very low frequency and eventually give room to the Debye scaling, that is allowed to appear only when $f_{u}=0$.
Figure 10: (a) The experimental density of states for liquid gallium at $500$ K, We thank Dehong Yu and the Authors of Ref. [52] to provide us the experimental data. (b) The INM DOS of Gallium simulated at different temperatures. The first-order melting temperature is between 320 K and 340 K. (c) The density of state function $S(\omega)$ derived from the Fourier transform of the VACF for the same simulated Gallium system.
In order to investigate this point, we have chosen Gallium as a benchmark example. As a reference, in Fig.10, we present the experimental data for the density of states at $500$ K from Ref.[52] together with our simulation results for the INM density of states $\rho(\omega)$ and the density of states function $S(\omega)$. Among the various features emerging from this comparison, we notice that at low temperature (e.g., $T=300$ K), $S(\omega)$ shows already a clear Debye scaling at low frequency and a vanishing diffusion constant ($S(0)=0$). On the contrary, the stable branch of the INM DOS is still strongly linear at low frequency, consistent with a large number of unstable mode still present at such a temperature. This difference is caused by the presence of non-diffusive unstable modes that survive in the solid phase. These modes are possibly a spurious effect of the INM computation and can be removed using various methods. We defer a detailed study of the nature of these modes for future work.
@@
In order to check this point further, in panel (b) of Fig.11, we have zoomed the stable part of the INM spectrum of Gallium at $T=1$ K and presented it in a log-log scale to highlight possible power-law scalings. The data present a clear linear scaling up to approximately $4.5$ meV, that is a consequence of the remaining unstable modes (and anharmonicities). On the other hand, around $\approx 4.5$ meV we observe a clear crossover to a larger power that is consistent with the Debye scaling $\rho(\omega)\propto\omega^{2}$ and that persists to $\approx 8$ meV. This reveals the approach of the solid-like behavior to low frequency by decreasing the temperature. By decreasing further $T$, one expects that the crossover frequency will move towards the origin, and finally vanish at $T=0$ when $f_{u}=0$. We notice that an alternative method to observed this trend, and the emergence of the Debye scaling, consists in nano-confining the liquid, as experimentally performed in [114]. We also notice that the presence of unstable modes in finite temperature (amorphous) solids has been verified experimentally, see for example [123].
@@ IV.5 Zero eigenvalue cusp singularity
Following previous observations [43, 44], Refs. [36, 45] have recently pointed out that the frequency representation of the INM density of states $\rho(\omega)$ might hide important information that are contained in the corresponding eigenvalue distribution $\upsilon(\lambda)$. Their arguments are supported by numerical simulations in a soft-sphere liquid model (data reproduced in panel (a) of Fig.14). Since $\lambda=\omega^{2}$, the eigenvalue distribution and the INM DOS are simply related via the Jacobian,
@@
Mathematically, it appears therefore unreasonable that one of the two quantities contain more information than the other. Nevertheless, in this Section we entertain this possibility and analyze in detail the arguments of [36, 45] in a large set of liquid models. First, we would like to point out that the statement that the INM density of states is universally linear at low frequencies is equivalent to the statement that $\rho(0)\neq 0$, that is certainly true. In other words, the presence of a cusp at $\lambda=0$ is not incompatible with a low-frequency form $\rho(\omega)=a(T)\omega+\dots$, observed in many instances.
@@
In order to characterize these features in a more quantitative way, we resort to the empirical function proposed in [45],
@@
Figure 12: The eigenvalue distribution $\upsilon(\lambda)$ for different temperatures. From top left to bottom right: Ar (40 bar) (a), Xe ($1.01$ bar) (b), Pb($1.01$ bar) (c), Ga($1.01$ bar) (d), N2 ($70.92$ bar) (e) and CS2 ($1.01$ bar) (f).
@@
In panel (a) of Fig.13, we show the temperature behavior of the fitted stable (red) and unstable (black) slopes as a function of temperature for the case of CS2. The slopes are extracted from the numerical data using Eq.(22). Similar results are obtained for the different liquids shown in Fig.12. We observe that the positive slope of the unstable branch grows with temperature while the negative one, corresponding to the unstable part of the spectrum, moves towards larger negative values by increasing $T$. This trend is opposite of what reported in [36, 45] for the soft sphere model. The reason for this difference is ultimately the opposite behavior of $\upsilon(0)$ with temperature in the soft sphere model. In particular, as evident from Fig.1 in [36], $\upsilon(0)$ in the soft sphere model decreases by increasing temperature. In our language, this corresponds to an INM slope $a(T)$ decreasing with temperature, as for LJ liquids, but opposite to our more realistic liquid systems. In fact, one could argue that the soft-sphere model is not a realistic model for liquids since it does not present a cohesive state. The relevance of the INM results in the soft-sphere model is therefore arguable, specially when compared to the more realistic cases considered in this work, that present opposite behavior.
@@
This contrast is then reflected in the temperature dependence of the difference of the slopes, that is presented in panel (b) of Fig.13. There, we observe that $\Delta_{\text{slope}}$ increases with temperature, following a trend approximately consistent with a linear power-law scaling. This is again very different from the $T^{-2/3}$ behavior observed in the soft sphere model [36, 45], not only for the precise power involved in the scaling, but even in the qualitative trend as well. Our findings imply that the $T^{-2/3}$ law found in [36, 45] is not universal and its value and significance should be revisited.
@@
Finally, we would like emphasize that the theoretical model presented in [36, 45] is not the only one able to predict the presence of a cusp-like singularity in the eigenvalue distribution of liquids. Indeed, a much older model known as the “soft potential model of liquids” (SPM) [30, 31] already captures such a feature. We provide a direct proof of this statement in panel b of Fig.14. The presence of a cusp singularity at $\lambda=0$ is evident. For completeness, in the inset we also show the INM DOS that shows the aforementioned linear in frequency regime despite the cusp.
@@
Figure 14: (a) The eigenvalue distribution $\upsilon(\lambda)$ for the soft-sphere liquid model presented in [45] and the corresponding fraction of unstable mode $f_{u}$ as a function of temperature in reduced LJ units. $T_{g}$ is the glass temperature of the corresponding system. (b) The eigenvalue distribution $\upsilon(\lambda)$ obtained from the theoretical soft potential model of liquids [30, 31] by converting the calculated $\rho(\omega)$ to $\upsilon(\lambda)$ with the Jacobian relation. The inset shows the original INM DOS $\rho(\omega)$ with the dashed black lines indicating the universal linear regime $\rho(\omega)\propto\omega$ which emerges clearly despite the presence of the cusp-like singularity in $\upsilon(\lambda)$.
@@
In this direction, it would be interesting to understand further the connection between the theory presented in [36, 45] and the older SPM [30, 31]. We notice that the two frameworks start from completely different assumptions. The heterogeneous elasticity theory of [36, 45] follows a harmonic description, whose validity is arguable in the liquid phase, especially at large temperatures. On the other hand, the SPM starts from a anharmonic picture of the dynamics that involves additional low-energy modes. Also in this case, the nature and origin of these modes in liquids remains an open question. String-like excitations [124] can represent a candidate for these anharmonic excitations.